OpenStudy (anonymous):

what are the zeros of the function f(x)=(5x^2-25x)/x

OpenStudy (elonasushchik):

In order to find the zeros, set the function equal to zero (y = 0). The given expression, f(x) = (5x^2 - 25x)/x can be simplified as follows: f(x) = x(5x - 25)/x then you can cancel the x from the numerator and the denominator which gives: f(x) =5x - 25 Subsequently, set f(x) = 0 5x - 25 = 0 5x = 25 x = 25/5 x =5 Hence, the only zero is 5 and not zero. I believe that the mistake you were making is not canceling the x's from the numerator and denominator..I assume you did the following: (5x^2 - 25x)/x = 0 5x^2 - 25x = 0 x (5x - 25) = 0 x = 0 and 5x - 25 = 0 5x = 25 x = 25/5 x = 5 which gave you two zeros: x = 0 and x = 5 So, the answer is just x = 5 Hope this helps!

OpenStudy (elonasushchik):

Medal?!

OpenStudy (anonymous):

thanks @ElonaSushchik idk how to give medal

OpenStudy (elonasushchik):

On the side it says best response. Click it and it will medal me:)

OpenStudy (elonasushchik):

Thanks!